OpenStudy (anonymous):

find the other coordinate in the point (5/6,Y) that lies on the unit circle how to get the Y if the X is 5/6 it should lies On the unit circle can i transpose it or X^2+Y^2=1?

OpenStudy (anonymous):

Y=1/2, i guess so.

OpenStudy (anonymous):

no not the answer

OpenStudy (anonymous):

ok give me a minute...

OpenStudy (anonymous):

ok.... (5/6)² + Y² = 1 25/36 + Y² = 1 Y² = 1 - 25/36 Y² = 11/36 Y = ±√11 / 6 <--- those are two distinct numbers. Hope you get this

OpenStudy (anonymous):

oh tnks a lot P.S sorry to dumb in math